Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.) 


μ = 28;  α = 3.6 


P(x ≥ 30) = _______ 
Answer #1

mean = 28

standard deviation = 3.6

P(x >= 30) = ?

consider,

P(x >= 30) = P((x-mu)/s >= (30 - 28)/3.6)

= P(z >= 0.5556)

= 0.2892422 [since from z table]

= 0.2892

P(x >= 30) = 0.2892